OpenStudy (anonymous):

Simplify the following product of complex numbers. (4 – 8i) * (-5 + 2i)

OpenStudy (ajprincess):

(4 – 8i) * (-5 + 2i) =-5(4 – 8i) + 2i(4 – 8i) =(-5*4)+(-5*-8i)+(2i*4)+(2i*-8i) =-20+40i+8i-16i^2 (i^2=-1) =-20+48i+(-16*-1) =-20+48i+16 =-4+48i

OpenStudy (anonymous):

gee got it 1/2 right, I will again go back and find out where i was off thanks for your help

OpenStudy (ajprincess):

yw.